### v1

# Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model

Abstract
Strongly interacting non-equilibrium systems are of great fundamental interest, yet their inherent complexity make then notoriously hard to analyze. We demonstrate that the dynamics of the Bose-Hubbard model, which by itself evades solvability, can be solved exactly at any interaction strength in the presence of loss tuned to a rate matching the hopping amplitude. Remarkably, the full solvability of the corresponding Liouvillian, and the integrability of the pertinent effective non-Hermitian Hamiltonian, survives the addition of disorder and generic boundary conditions. By analyzing the Bethe ansatz solutions we find that even weak interactions change the qualitative features of the system, leading to an intricate dynamical phase diagram featuring non-Hermitian Mott-skin effects, disorder induced localization, highly degenerate exceptional points, and a Bose glass-like phase of fragmented condensates. We discuss realistic implementations of this model with cold atoms.
Introduction.– In cold atom systems with strong interactions [RevModPhys.80.885, Kinoshita2006, Lewenstein2007, Gross2017, Zhang2018] and disorder [Garreau2017], dissipation is an ever-present adversary, since environmental coupling tends to destroy intriguing, but delicate, quantum phenomena. However, it is being increasingly appreciated that dissipation itself may in fact be harnessed [diehl2008quantum, Bardyn_2013, goldman2016topological, Müller2021, Helmrich2020, Sponselee_2019, Gou_2020, Liang_2022, Bouchoule2021, Yamamoto2023, mengdissipative, fandissipative, Gong2018, ashida2020non, bergholtzreview] and induce unique intriguing effects [Gong2018, ashida2020non, bergholtzreview], thus providing a new frontier in the study of non-equilibrium quantum many-body systems [eisert2015quantum]. The vast majority of recent studies have focused on non-interacting dissipative systems [Gong2018, ashida2020non, bergholtzreview]. In the quantum realm these systems may realize novel topological phenomena such as the Liouvillian version [Song2019, PhysRevLett.127.070402, PhysRevResearch.4.023160] of the non-Hermitian skin effect [lee2016, yao2018, kunst2018, martinez2018, Okuma2020, Okuma2023, lin2023] in which a macroscopic number of eigenstates are exponentially localized at the boundaries of the system. Several recent studies investigate the interplay between interactions and the skin effect [Yoshida2023, Kim2023, Yang2021, Mao_2023, Wang_2023, zheng2023exact, fate22, Shen_2022, hamanaka2024multifractality, zhong2024density], yet this topic remains far less explored than the non-interacting case.
Even in absence of dissipation, strongly interacting models are difficult to analyze. This in particular applies to the paradigmatic one-dimensional Bose-Hubbard model [bosehubbard, bosehubbardtool], which is famously non-integrable [Choy1982] for more than two sites [Links2003], in contrast to its fermionic counterpart. The absence of integrability combined with its key experimental importance [Jaksch1998] has made it a popular target in numerical studies, using, for example density, renormalization group techniques [Kollath_2004, Kuehner_1998, Ejima_2011, Urba_2006, Schmidt_2007, kollath2007], Monte Carlo methods [Kashurnikov_1996, Pollet_2013], and exact diagonalization [Sowinski_2012, Zhang_2010, kollath2007].
A seminal exact solution strategy of a strongly interacting many body system was provided by Bethe [Bethe] in his famous exact solution of the Heisenberg model. Much more recently, third quantization [Prosen_2008] was invented as a framework to solve quadratic (non-interacting) Liouvillians, describing driven-dissipative systems, and in the last few years the notion of integrability was extended to more general systems including certain interacting Liouvillians [Medvedyeva_2016, de_Leeuw_2021, Ziolkowska_2020, Buča2020, Nakagawa2021]. The application of integrability methods to the solution of Liovillians has thus made it possible to construct exactly solvable models that combine both dissipation and strong interactions, but exactly solvable models of that kind are rare, and exactly solvable models that incorporate disorder, dissipation and interactions simultaneously remain elusive.
In this letter, we present and analytically solve a disordered Bose-Hubbard chain with environmental coupling that is tailored to render it exactly solvable while remaining realistic in cold atom systems (see Fig. 1). In the process we derive and solve the non-Hermitian Hamiltonian of the unidirectional Bose-Hubbard chain that was first shown to be Yang-Baxter integrable in Ref. zheng2023exact. Here we extend this analysis in several ways that reveal qualitatively new physics. We show how the model emerges as an effective description of a realistic setup, and that it remains integrable in the presence of arbitrary on-site potentials, including random on-site disorder, and with open boundary conditions. The open boundary conditions enable non-Hermitian Mott-skin effects and highly degenerate exceptional points, and adding disorder leads to a novel phase that is reminiscent of a Bose glass. Furthermore, we stress that our analysis goes beyond the effective non-Hermitian Hamiltonian approach by providing a complete solution of the Lindblad Liouvillian.
Model and physical realization.–
Figure 1: The models under consideration in this letter. (a) A triangular ladder as implemented in Refs. Gou_2020 and Liang_2022, with an additional on-site potential and interactions of strength $U$. The upper sites have dissipation with rate $\Gamma$, while the lower ones do not. A fictitious magnetic flux $\phi$ leads to a relative phase in the hopping terms between the different sites Upon integrating out the green sites we obtain (b) with dissipation acting on neighbouring sites, where the magnetic flux has been chosen to be $\phi=-\frac{\pi}{2}$. (c) Shows the system described by the integrable unidirectional effective non-Hermitian (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model), where the dissipation has been absorbed into the hopping parameter and the on-site potential.
Dissipative cold atom systems may be realistically modeled using the Lindblad master equation [Lindblad, Breuer],
$\frac{d\rho}{dt}=\mathcal{L}\rho=-i[H,\rho]+\sum_{j}\gamma_{j}\left(2\ell_{j}\rho\ell^{\dagger}_{j}-\{\ell_{j}\ell_{j}^{\dagger},\rho\}\right).$ (1)
with $\gamma_{j}\geq 0$. We consider the coherent dynamics to be described by the disordered Bose-Hubbard Hamiltonian
$H=-$ $t\sum_{j=1}^{N}a_{j}^{\dagger}a_{j+1}+a_{j+1}^{\dagger}a_{j}+\frac{U}{2}\sum_{j=1}^{N}n_{j}(n_{j}-1)-\sum_{j=1}^{N}\mu_{j}n_{j},$
$-\epsilon(a^{\dagger}_{N}a_{1}+a^{\dagger}_{1}a_{N})$ (2)
where $a_{j}$ are bosonic annihilation operators, $n_{j}=a^{\dagger}_{j}a_{j}$ are the number operators, and $\mu_{j}$ are random disorder variables.
One way to incorporate realistic dissipation so that the system becomes exactly solvable is to consider a triangular ladder [Liang_2022, Gou_2020], as in Fig. 1(a). Particles dissipate from the sites on the top and these sites can be integrated out [supplemental_material] (see also Ref. Kamenevbook, leaving a chain with jump operators
$\ell_{j}=a_{j}+ia_{j+1}$ (3)
as in Fig. 1(b), cf. Ref. Song2019. The relative phase factor may be obtained experimentally by including a magnetic field. Remarkably, at the level of the effective non-Hermitian Hamiltonian, the hopping in one direction vanishes if $\gamma_{j}=t$ for $j\neq N$ and $\gamma_{N}=\frac{\epsilon}{2}$. Explicitly, the resulting effective non-Hermitian Hamiltonian (Fig. 1(c)) becomes
$H_{eff}$ $=-2t\sum_{j=1}^{N}a_{j+1}^{\dagger}a_{j}+\frac{U}{2}\sum_{j=1}^{N}n_{j}(n_{j}-1)-\sum_{j=1}^{N}\mu_{j}^{\prime}n_{j}$
$-\epsilon a^{\dagger}_{1}a_{N}.$ (4)
where we defined $\mu^{\prime}_{j}=\mu_{j}-2it$ for $j\neq 1,N$ and $\mu^{\prime}_{1}=\mu_{1}-i(t+\frac{\epsilon}{2}),\mu^{\prime}_{N}=\mu_{N}-i(t+\frac{\epsilon}{2})$. Ref. Liang_2022 realized and analyzed the non-interacting and non-disordered limit of Eq. (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model) in a cold atom system. Here we include both interactions and disorder, and in addition to Eq. (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model), also consider the full Liovillian quantum dynamics as described by Eqs. (1), (2) and (3).
Bethe ansatz solution.– Using the machinery of the quantum inverse scattering method [Korepinbook] it may be shown that the right eigenvectors of the Hamiltonian, and of all its constants of motion defined in the supplemental material [supplemental_material], have the form given by the algebraic Bethe Ansatz
|$\ket{\{\lambda_{n}\}}=\prod_{n=1}^{M}\mathcal{B}(\lambda_{n})\ket{\Omega}$ (5)
[…]
$\prod_{k=1}^{N}(\lambda_{n}-\mu^{\prime}_{k})$ $\prod_{j\neq n}(\lambda_{n}-\lambda_{j}-U)$
$-\epsilon\left(-\frac{2t}{U}\right)^{N}\prod_{j\neq n}(\lambda_{j}-\lambda_{n}-U)=0.$ (6)
In terms of the rapidities the eigenvalue corresponding to the eigenstate (5) of the non-Hermitian effective Hamiltonian is $E(\{\lambda_{n}\})=\sum_{n=1}^{M}\lambda_{n}$. In the case without an onsite potential, meaning $\mu_{k}^{\prime}=0$ for all $k$, and with $\epsilon\neq 0,$ this was first shown in Ref. zheng2023exact.
The coordinate Bethe Ansatz in terms of the rapidities is [Jiang_2020]
$\Psi\left(\{x_{k}\}_{k=1}^{M}\right)=\sum_{\sigma\in S_{M}}\prod_{j>k}f(\lambda_{\sigma_{j}}-\lambda_{\sigma_{k}})\prod_{n=1}^{M}\prod_{j=1}^{x_{n}}\left(\frac{\lambda_{\sigma_{n}}-\mu_{j}}{-2t}\right)$ (7)
where $f(\lambda)=1-\frac{U}{\lambda}$ and $S_{M}$ is the set of all permutations of the integers between $1$ and $M$. This expression matches the algebraic Bethe Ansatz.
Integrability and solvability.– The integrability structure outlined above can be used to explicitly diagonalize the Liouvillian (1). It is important that there is only dissipation (or only gain), and that the effective non-Hermitian Hamiltonian commutes with the number operator, as this ensures that the Liouvillian is triangular in the basis given by states of the form |$\ket{\{\lambda_{n}\}_{n=1}^{M}}\bra{\{\nu_{m}\}_{m=1}^{M^{\prime}}}$ | [Torres2014]. Its eigenvalues are given by
$\omega(\{\lambda_{n}\}_{n=1}^{M},\{\nu_{m}\}_{m=1}^{M^{\prime}})=-i\sum_{n=1}^{M}\lambda_{n}+i\sum_{m=1}^{M^{\prime}}\bar{\nu}_{m}$ (8)
and the eigenoperators have the form
$\rho_{(\lambda,\nu)}=\sum_{n=0}^{m}\sum_{(\lambda^{\prime},\nu^{\prime})}v^{(l,m;n)}_{(\lambda,\nu);(\lambda^{\prime},\nu^{\prime})}\ket{\{\lambda^{\prime}\}}\bra{\{\nu^{\prime}\}},$ (9)
where the coefficients are given in the supplemental material [supplemental_material] (see also Refs. Slavnov1989, Gaudin, Oota2004, Brody2014, Slavnovbook). Note that the unique steady state of the system is the empty state, so there cannot be any conserved quantities. The Liouvillian therefore cannot easily be said to be integrable, in spite of the underlying integrability structure that allows its exact analytical solution.
Correlation functions and phase diagram for open boundary conditions.– We now examine the properties of the model defined by the effective Hamiltonian (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model) with random disorder and open boundary conditions ($\epsilon=0$), which exhibits a remarkable degree of solvability. As we detail below this leads to the phase diagram in Fig. 2. For arbitrary disorder the Bethe equations now take the factorized form
$\prod_{k=1}^{N}(\lambda_{n}-\mu_{k}^{\prime})\prod_{j\neq n}(\lambda_{n}-\lambda_{j}-U)=0$ (10)
which have solutions $\lambda_{n}^{k}=\mu_{k}^{\prime}+(n-1)U$ where $n=1,2,\ldots$. Inserting this into the expression for the energy we see that the spectrum is given by
$E(n_{1},n_{2},\ldots,n_{N})=\sum_{k=1}^{N}\mu_{k}^{\prime}n_{k}+\frac{U}{2}n_{k}(n_{k}-1).$ (11)
The effective Hamiltonian exhibits exceptional points, i.e. at a non-Hermitian degeneracy at which both eigenvalues and eigenstates coalesce, whenever there are distinct partitions of integers, $\{n_{i}\}$, giving identical energies in Eq. (11). In particular this happens when any two (or more) disorder variables, $\mu_{i}$, are equal. Moreover, if the potential is constant (no disorder), there are exceptional points of very high order. It is worth noting that although at $t=0$ the integers $\{n_{i}\}$ labeling energies and eigenstates correspond to particle densities, their interpretation is less straightforward at $t\neq 0$.
Figure 2: The ground state phase diagram of the non-Hermitian Hamiltonian with open boundary conditions at integer filling. For $U>2\Delta$ the ground state is in a Mott phase, which can either be localized to one of the boundaries or to the right depending on $t$, whereas for $U<2\Delta$ the system is in Bose-Glass-like phase with several different condensates, which may also be localized to either the left or the right. At $U=0$ there is a transition between a Bose-Einstein condensate at the boundary and a Bose-Einstein condensate at a single site in the bulk.
In the limit $U\rightarrow 0$ it is sufficient to treat the single-particle eigenstates. These are given by
$\Psi_{\lambda}(x)=\prod_{j=1}^{x}\left(\frac{\lambda-\mu_{j}^{\prime}}{-2t}\right)$ (12)
which are eigenstates if $\lambda=\mu_{n}^{\prime}$ for $n=1,2,\ldots N$. If there is no disorder, meaning $\mu_{n}^{\prime}=0,$ all the eigenstates coalesce to $\psi(x)=\delta_{x,1}$. One simple observation is in that if $\lambda=\mu_{n}^{\prime}$ then the particle density is 00 for $x>m$ where $m$ is the smallest integer such that $\mu_{m}^{\prime}=\mu_{n}^{\prime}$. The particle is therefore sharply localized to a finite region. In addition, if $t>>|\mu_{m^{\prime}}^{\prime}-\mu_{n^{\prime}}^{\prime}|$ for all $m^{\prime},n^{\prime}$ we see that the particle is exponentially localized in the left side of that region for all eigenstates. On the other hand, if $t<<|\mu_{m^{\prime}}^{\prime}-\mu_{n^{\prime}}^{\prime}|$ on average each eigenstate is exponentially localized to its corresponding site. Extending to the many-body ground state, we see that it is a gapless condensate of particles in the state with $\lambda$ equal to the $\mu_{n}^{\prime}$ with smallest real part. Due to the simplicity of this state, it is possible to obtain the asymptotics of the two-point correlation function. Let $x_{2}>x_{1}$ both be in the region where the wave function is non-zero. Then, using the notation $\mu_{k}=\mu_{k}^{\prime}+2it$
$\bra{GS}a^{\dagger}_{x_{1}}a_{x_{2}}\ket{GS}\sim\prod_{n=1}^{x_{1}}\left(\frac{\mu_{n}-\mu_{x}}{-2t}\right)^{2}\prod_{n=x_{1}+1}^{x_{2}}\left(\frac{\mu_{n}-\mu_{x}}{-2t}\right).$ (13)
In order to proceed we specify that the disorder variables $\mu_{k}$ are uniformly distributed in the interval $[-\Delta,\Delta]$. Since this is the ground state $\mu_{n}$ must be the smallest disorder variable, which is $-\Delta$ in the thermodynamic limit. Including the wave function normalization and averaging over the disorder we find that the asymptotics of the two point correlator is
$G(x_{1},x_{2})$ $\propto\left(\frac{\Delta}{-t}\right)^{2x_{1}}\left(\frac{\Delta}{-2t}\right)^{x_{2}-x_{1}}.$ (14)
As expected this is not translationally invariant, and we see that the correlation length $\xi=1/\ln\frac{\Delta}{2t}$ which diverges at $\Delta=2t,$ signifying a phase transition. One of the phases is a skin-effect-phase, where all the particles are localized at one of the edges (corresponding to the red line at the left edge of Fig. 2), and the other phase is a disorder-induced localized phase, where all the particles are localized at some site (corresponding to the green line at the left edge of Fig. 2). Other states show a similar behaviour.
Now we let $U>0$. This has dramatic effects for arbitrarily weak interactions in the thermodynamic limit, $N\rightarrow\infty$. In the ground state the particles are no longer condensed into the same mode, due to the repulsive interactions. The particles spread out over the different modes, so that there are competing condensates. This is reminiscent of a Bose-Glass [Fisher1989, Scalettar1991], where the competition between disorder and interactions causes disjoint localized pockets of superfluid condensate to form. Here, in contrast, the condensates are not spatially disjoint unless the disorder is much larger than the hopping parameter, as in the non-interacting case.
If $U$ is sufficiently small one of the modes will still have a most of the particles, say $M_{0}$, so that there are rapidities $\lambda_{n}=-\Delta+(n-1)U$ for $n=1,2,\ldots M_{0}$. The magnitude of the disorder part in the position dependent factor in the wave function is always smaller than $2\Delta,$ so a factor $\left(\frac{(M_{0}-1)U+2\Delta}{-2t}\right)^{x_{1}+x_{2}}$ may be extracted from the sum, so that for $x_{2}>>x_{1}$ the leading behaviour of the two-point correlator is approximately
$G(x_{1},x_{2})\sim\left(C\frac{(M_{0}-1)U+2\Delta}{-2t}\right)^{x_{2}-x_{1}}$ (15)
for some real number $C\sim 1$ which arises from the disorder average. We see that the correlation length is $\xi=1/\ln\left(\frac{(M_{0}-1)U+2\Delta}{2t}\right).$ Its divergence at $2t=2\Delta+(M_{0}-1)U$ marks a transition between a boundary-localized (the purple part of Fig. 2) and a site-localized (the green part of Fig. 2) phase.
If $U>2\Delta$ the energy is minimized by equally distributing the integers in Eq. (11). Suppose for simplicity that $M/N=n$ is an integer. In that case the gap is $U-2\Delta$, so this ground state is a Mott insulator at integer filling. The leading order behaviour of the two-point correlator for $x_{2}>>x_{1}$ is
$G(x_{1},x_{2})\sim\left(C\frac{(n-1)U+2\Delta}{-2t}\right)^{x_{2}-x_{1}}$ (16)
again for some real number $C\sim 1$. Note that at $n=1$ the leading order behaviour is independent of $U$ (up to the constant $C,$ which may have some $U$ dependence). There is a phase transition at $(n-1)U+2\Delta=\frac{2t}{C}$ between a site-localized and a boundary-localized Mott insulator, as shown in the yellow and red part of Fig. 2 respectively.
Liouvillian dynamics.– Now let us turn to the dynamics of the full Lindblad master equation. For a generic initial state we expect that the time evolution includes every eigenoperator (9), hence the two-point correlator consists of some time-dependent linear combination of all two-point correlators of the model defined by the Hamiltonian (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model). By the same reasoning as for the ground state correlators the leading order time dependent asymptotics are
$G(x_{1},x_{2};\tau)=\sum_{n=1}^{M}D_{n}(\tau)\left(C_{n}\frac{(n-1)U+2\Delta}{-2t}\right)^{x_{2}-x_{1}},$ (17)
where $\tau$ is the time, $C_{n}\sim 1$ are constants, and $D_{n}(\tau)$ are obtained from the coefficients in Eq. (9). Since there is only dissipation it follows that the functions $D_{n}(\tau)\rightarrow 0$ as $\tau\rightarrow\infty$. We see that if $U$ is non-zero there will only be a phase transition between boundary localization and site localization once the mean particle number has decreased to a finite number. If $U=0,$ on the other hand, there is always such a dynamic phase transition for $t=\Delta$.
Discussion.– In this work, we have shown that the paradigmatic example of a non-integrable model, namely the Bose-Hubbard model, remarkably becomes exactly Bethe Ansatz solvable in the presence of dissipation terms that may be realistically implemented with cold atoms. The effective non-Hermitian Hamiltonian as well as the full non-equilibrium quantum dynamics feature a rich phase diagram that invites ample future theoretical as well as experimental work.
As mentioned above, the realization of a momentum-space lattice version of this model in the non-interacting limit in an cold atom setup was reported in Ref. Gou_2020. There the parameters were chosen so that interactions were unimportant and attractive. We expect that similar setups can be tuned to the regime of repulsive interactions described in this paper, although the real-space interactions must be chosen to be attractive in order the be repulsive in the momentum-space lattice [An_2018]. While previous realizations of the non-interacting limit have used ${}^{87}\mathrm{Rb}$ atoms, ${}^{39}\mathrm{K}$ may be more suitable for the interacting system, as its relevant Feschbach resonances are wider, allowing for more precise control of the interactions [Meierthesis].
The model studied in this work may be generalized to higher dimensions and more involved interaction terms. To this end one has to devise Lindblad operators such that the particles in the system can only hop in one direction along each dimension. Then the Fock basis can be ordered so that the kinetic term is triangular, and consequently the non-hermitian effective Hamiltonian and the Liouvillian are triangular as long as the remaining terms are polynomial in the on-site densities. If open boundary conditions are chosen it is straightforward to calculate the eigenvalues of the effective Hamiltonian, while determining the corresponding eigenvectors is technically more complicated.
An interesting conceptual aspect of this work is that the Lindbladian does not have any conserved quantities, yet it is solvable by the Bethe Ansatz, similarly to the models discussed in [Buča2020, Nakagawa2021]. There, however, the system is integrable in the absence of dissipation in contrast to the present case [Choy1982].
The topological nature of the non-interacting non-Hermitian skin effect suggests that the quantum many-body generalizations thereof studied here is also robust to generic perturbations, which is promising in the context of its realization.
Finally, we also note that an alternative interpretation of the non-Hermitian Hamiltonian Eq.  (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model) can be gleaned from [Hatano1996, Hatano1997, Lehrer1998]. There the difference in the hopping to the left and to the right, respectively, are obtained from the strength of a magnetic field applied to interacting flux lines in an array of defects in a superconductor. The unidirectional Hamiltonian is obtained when the transverse field component is large.
Acknowledgements.
Acknowledgements.– We acknowledge helpful discussions with Rodrigo Arouca de Albuquerque and Alexander Fagerlund. This work was supported by the Swedish Research Council (VR, grant 2018-00313), the Wallenberg Academy Fellows program of the Knut and Alice Wallenberg Foundation (2018.0460) and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine.